A total of $9000 is invested in two funds paying 5% and 6% annual interest. The combined interst is $510. How much of the $9000 is invested in each fund?
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

.06X+.05(9000-X)=510
.06X+450-.05X=510
.01X=510-450
.01X=60
X=60/.01
X=6000 INVESTED @ 6%
9000-6000=3000 INVESTED @ 5%
PROOF
.06*6000+.05*3000=510
360+150=510
510=510

Answer by stanbon(75887)

A total of $9000 is invested in two funds paying 5% and 6% annual interest. The combined interst is $510. How much of the $9000 is invested in each fund?
-----------
Let amt. invested at 5% be "x" ; Interest on this is 0.05x
Amt. invested at 6% is "9000-x" ; Intereston this in 0.06(9000-x)= 540-0.06x
----------
EQUATION:
interest + interest = 510
0.05x+540-0.06x = 510
-0.01x = -30
x=$3000 (amt. invested at 5%)
9000-3000=$6000 (amt. invested at 6%)
--------
Cheers,
Stan H.